2. Two projects being considered are mutually exclusive and have the following projected cash flows:. If the
required rate of return on these projects is 11 percent, which would be chosen and why?
Project A Project B
Year Cash Flow Cash Flow
0 -40,000 -40,000
1 15,625 0
2 15,625 0
3 15,625 0
4 15,625 0
5 15,625 99,500
3. Two projects being considered are mutually exclusive and have the following projected cash flows:. If the
required rate of return on these projects is 6 percent, what is each projects NPV?
Project A Project B
Year Cash Flow Cash Flow
0 -80,000 -80,000
1 20,000 0
2 20,000 0
3 20,000 0
4 20,000 0
5 20,000 100,000
Answer #1

2.

H I J K
Project A
6 Discount rate 0.11
7 Year Cash flows PVF@11% = 1 (1+r)^n Present Value
8 0 -40000 1.000000 -22000.00
9 1 15625 0.900901 14076.58
10 2 15625 0.811622 12681.60
11 3 15625 0.731191 11424.87
12 4 15625 0.658731 10292.67
13 5 15625 0.593451 9272.68
14
15 NPV 35748.39
16
PROJECT B
H I J K
19 Project B
20 Discount rate 0.11
21 Year Cash flows PVF@11% = 1 (1+r)^n Present Value
22 0 -40000 1.000000 -22000.00
23 1 0 0.900901 0.00
24 2 0 0.811622 0.00
25 3 0 0.731191 0.00
26 4 0 0.658731 0.00
27 5 99500 0.593451 59048.41
28
29 NPV 37048.41

FORMULA:

H I J K
Project A
6 Discount rate 0.11
7 Year Cash flows PVF@11% = 1 (1+r)^n Present Value
8 0 -40000 =1/(1+$K$6)^H8 -22000
9 1 15625 =1/(1+$K$6)^H9 =I9*J9
10 2 15625 =1/(1+$K$6)^H10 =I10*J10
11 3 15625 =1/(1+$K$6)^H11 =I11*J11
12 4 15625 =1/(1+$K$6)^H12 =I12*J12
13 5 15625 =1/(1+$K$6)^H13 =I13*J13
14
15 NPV =SUM(K8:K13)
16
PROJECT B
H I J K
19 Project B
20 Discount rate 0.11
21 Year Cash flows PVF@11% = 1 (1+r)^n Present Value
22 0 -40000 =1/(1+$K$6)^H22 -22000
23 1 0 =1/(1+$K$6)^H23 =I23*J23
24 2 0 =1/(1+$K$6)^H24 =I24*J24
25 3 0 =1/(1+$K$6)^H25 =I25*J25
26 4 0 =1/(1+$K$6)^H26 =I26*J26
27 5 99500 =1/(1+$K$6)^H27 =I27*J27
28
29 NPV =SUM(K22:K27)
PROJECT A PROJECT B
IRR IRR
A B C D
Year Cash flow Year Cash flow
1 0 -40000 1 0 -40000
2 1 15625 2 1 0
3 2 15625 3 2 0
4 3 15625 4 3 0
5 4 15625 5 4 0
6 5 15625 6 5 99500
7 7
8 =IRR(B1:B6) 8 =IRR(D1:D6)
9 IRR 27.44% 9 IRR 19.99%

FORMULA:

PROJECT A PROJECT B
IRR IRR
A B C D
Year Cash flow Year Cash flow
1 0 -40000 1 0 -40000
2 1 15625 2 1 0
3 2 15625 3 2 0
4 3 15625 4 3 0
5 4 15625 5 4 0
6 5 15625 6 5 99500
7 7
8 =IRR(B1:B6) 8 =IRR(D1:D6)
9 IRR =IRR(O9:O15) 9 IRR =IRR(S9:S15)

PROJECT B HAS HIGHEST NPV SO ITS CHOOSEN BY COMPARING NPV AND IRR.

3.

H I J K
Project A
6 Discount rate 0.06
7 Year Cash flows PVF@6% = 1 (1+r)^n Present Value
8 0 -80000 1.000000 -22000.00
9 1 20000 0.943396 18867.92
10 2 20000 0.889996 17799.93
11 3 20000 0.839619 16792.39
12 4 20000 0.792094 15841.87
13 5 20000 0.747258 14945.16
14
15 NPV 62247.28
16
PROJECT B
H I J K
19 Project B
20 Discount rate 0.06
21 Year Cash flows PVF@6% = 1 (1+r)^n Present Value
22 0 -80000 1.000000 -22000.00
23 1 0 0.943396 0.00
24 2 0 0.889996 0.00
25 3 0 0.839619 0.00
26 4 0 0.792094 0.00
27 5 100000 0.747258 74725.82
28
29 NPV 52725.82

PROJECT 'A' HAS HIGHEST NPV THAN PROJECT B.

FORMULA:

H I J K
Project A
6 Discount rate 0.06
7 Year Cash flows PVF@6% = 1 (1+r)^n Present Value
8 0 -80000 =1/(1+$K$6)^H8 -22000
9 1 20000 =1/(1+$K$6)^H9 =I9*J9
10 2 20000 =1/(1+$K$6)^H10 =I10*J10
11 3 20000 =1/(1+$K$6)^H11 =I11*J11
12 4 20000 =1/(1+$K$6)^H12 =I12*J12
13 5 20000 =1/(1+$K$6)^H13 =I13*J13
14
15 NPV =SUM(K8:K13)
16
PROJECT B
H I J K
19 Project B
20 Discount rate 0.06
21 Year Cash flows PVF@6% = 1 (1+r)^n Present Value
22 0 -80000 =1/(1+$K$6)^H22 -22000
23 1 0 =1/(1+$K$6)^H23 =I23*J23
24 2 0 =1/(1+$K$6)^H24 =I24*J24
25 3 0 =1/(1+$K$6)^H25 =I25*J25
26 4 0 =1/(1+$K$6)^H26 =I26*J26
27 5 100000 =1/(1+$K$6)^H27 =I27*J27
28
29 NPV =SUM(K22:K27)